OpenStudy (anonymous):

The length of a side and a diagonal of a rhombus are 5cm and 8cm respectively. Find the area of the rhombus?

OpenStudy (agent0smith):



OpenStudy (mathstudent55):

 In a rhombus, the diagonals are perpendicular bisectors.

OpenStudy (anonymous):

As we know that the halves of diagonals and a side of a rhombus from a right angled triangle with side as hypotenuse. Let the length of the other diagonal =2x. ⇒x^2+(8/2)2= 52 ⇒x^2+16=25 ⇒x^2=25-16 ⇒x^2=9 ⇒x=3 The measure of the another diagonal of the rhombus = 2×3 = 6 cm ∴Area of the rhombus = (1/2)×Diagonal of a rhombus×Measure of the another diagonal ⇒Area of the rhombus = (1/2) ×8×6=24 cm^2

OpenStudy (anonymous):

Did i do the sum correct or not.

OpenStudy (agent0smith):

Be careful how you write in your equations... you have 5^2 written as 52, and (8/2)^2 written as (8/2)2. But yes, your working looks correct.

OpenStudy (anonymous):

yup..u r write...